Determine the pH of a 0.125 M Na2CO3 solution. (Ka of H2CO3 = 5.6 x 10-11)
Answer #1

Sol . As Initial Concentration of Na2CO3 = [Na2CO3]   = 0.125 M

So , [ Na2CO3 ] = [ CO32- ] = 0.125 M  

Now ,

Reaction of CO32- with H2O :

CO32- + H2O <----> HCO3-   + OH-

Initial 0.125 0 0

Change - x + x + x

Equilibrium 0.125 - x x x

Also , Kb = Kw / Ka = 10-14 / ( 5.6 × 10-11 )

= 0.1785 × 10-3

So , Kb = [HCO3-] [OH-] / [CO32-]

0.1785 × 10-3 = x2 / ( 0.125 - x )

As Kb is very small , so , 0.125 - x = 0.125  

Therefore , 0.1785 × 10-3 × 0.125 = x2

x2 = 0.223125 × 10-4

x = 0.4724 × 10-2

So , [OH-] = x = 0.4724 × 10-2 M  

Now , pOH = - log[OH-] = - log (0.4724 × 10-2 )  

= 2.32

Therefore , pH = 14 - pOH = 14 - 2.32 = 11.68